# Quantum dispersionless KdV hierarchy revisited

### v2

@@ Acknowledgement
The author is a JSPS International Research Fellow and his research is supported by JSPS KAKENHI Grant Number 23KF0114. He would like to thank Wenda Fang, Si-Qi Liu, Youjin Zhang and Chunhui Zhou for very helpful discussions.
@@ 2.1. Basics of vertex algebra
Let us review some basic facts about vertex algebras, and we refer to [8] for details. The paper [4] also serves as a good reference for the topic.
Recall that a vertex algebra is given by the data $\left(V,|0\rangle,T,Y(-,z)\right)$, where $V$ is a vector space over $\mathbb{C}$ called the space of states, $|0\rangle\in V$ is a special vector called the vacuum vector, $T\in\mathrm{End}(V)$ is a linear operator called the infinitesimal translation operator and $Y(-,z)\colon V\to\mathrm{End}(V)[[z^{\pm 1}]]$ is a linear morphism called the state-field correspondence. For each state $a\in V$, let us denote
@@
We call $a_{(n)}$ the $n$-th product by $a$. These data are required to satisfy the following axioms:
@@
  • $T|0\rangle=0$ and for any $a\in V$, $[T,Y(a,z)]=\partial_{z}Y(a,z)$.
@@
$[a_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}a_{(n)}b.$
The $\lambda$-bracket satisfies the following properties, where many useful identities for vertex algebras are encoded:
@@
    $[Ta_{\lambda}b]=-\lambda[a_{\lambda}b]$, and $[a_{\lambda}Tb]=(T+\lambda)[a_{\lambda}b]$.
@@
    $[a_{\lambda}[b_{\mu}c]]=[b_{\mu}[a_{\lambda}c]]+[[a_{\lambda}b]_{\lambda+\mu}c]$.
[…]
@@
[…]
@@
(2.3) $[a_{\lambda}:bc:]=:[a_{\lambda}b]c:+:b[a_{\lambda}c]:+\int_{0}^{\lambda}[[a_{\lambda}b]_{\mu}c]d\mu.$
Another standard fact about the normal product is that
@@
In this paper, we will only consider the Heisenberg vertex algebra. Let $V$ be the Heisenberg vertex algebra of rank $N$ generated by states $b^{1},\dots,b^{N}$ satisfying the commutation relation
(2.5) $\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N,$
where $(\eta^{\alpha\beta})$ is a constant non-degenerate symmetric matrix and $\hbar$ is just a formal parameter. The state space $V$ is a free $\mathbb{C}[\hbar]$-module spanned by states of the form
@@
The action of each $b^{\alpha}_{(n)}$ on state vectors is then uniquely determined by the relation (2.5). Note that in particular we have $b^{\alpha}_{(0)}=0$. Note that it follows from the vacuum axiom of $V$ that $b^{\alpha}=b^{\alpha}_{(-1)}|0\rangle$. For the Heisenberg vertex algebra, the state-field correspondence takes a very specific form, which can be described as follows.
@@
$Y\left(a,z\right)=(-1)^{\sum k_{i}}\sum_{m_{1},\dots,m_{n}}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:z^{-n-\sum m_{i}},$
where the normal order on the right-hand side means that we always put annihilators, namely those operators $b^{\alpha}_{(m)}$ with $m\geq 0$, on the right.
Remark 2.1.
Note that the normal order defined above is well-defined due to the commutation relation (2.5). For general vertex algebras, we cannot define such normal order.
@@
Let us prove by induction on $n$. The case $n=0$ is trivial. Assume that we have proved the proposition for some $n\geq 0$, then let us consider the state
$a=b^{\alpha}_{(-k-1)}b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V.$
First we notice that
@@
Then by using the fact that $Y(T^{k}b^{\alpha},z)=\partial_{z}^{k}Y(b^{\alpha},z)$ and the identity (2.4), we arrive at the following expression for $Y(a,z)$:
$\begin{split}(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\\ m<0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot b^{\alpha}_{(m-k)}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:z^{-n-\sum m_{i}-m-1}\end{split}$
$\begin{split}+(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\\ m\geq 0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:b^{\alpha}_{(m-k)}z^{-n-\sum m_{i}-m-1}.\end{split}$
In the first summation, note that $m-k<0$, hence by definition of the normal order, we have
$b^{\alpha}_{(m-k)}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\,=\,:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:,\quad m<0.$
In the second summation, note that for $10\leq m\leq k-$ 1, the binomial coefficient $\binom{m}{k}=0$, and for $m\geq k$ we have
$:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:b^{\alpha}_{(m-k)}=:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:.$
@@
Fix a state $a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$ for some $n\geq 0$ and $k_{i}\geq 0$, then we have
$a_{(m)}=\sum_{\begin{subarray}{c}j_{1}\leq\dots\leq j_{n}\\ \beta_{1},\dots,\beta_{n}\end{subarray}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})}\in\mathrm{End}(V),$
here $C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}$ are some constants (depending on the state $a$ and $m\in\mathbb{Z}$).
@@ 2.2. Quantum Hamiltonian structure
In this section, we construct a deformation quantization of Hamiltonian structure of hydrodynamic types using the Heisenberg vertex algebra. To this end, we will first recall the notion of a (classical) Hamiltonian structure, and one may refer to [16] for an introduction to the general theory of Hamiltonian structures.
@@
$\mathcal{A}=\mathbb{C}\left[v^{\alpha,p}\colon\alpha=1,\dots,N,\,p\geq 0\right].$
Define a derivation $\partial_{x}$ of $\mathcal{A}$ by
@@
Let us consider the following special Hamiltonian structure of hydrodynamic type:
@@
where $(\eta^{\alpha\beta})$ is a non-degenerate symmetric constant matrix. It is proved in [9] that the bracket (2.6) serves as a universal local model for Hamiltonian structures with hydrodynamic leading terms. Such Hamiltonian structures are ubiquitous in the theory of integrable hierarchies. For example, all Dubrovin-Zhang hierarchies possess this kind of Hamiltonian structures.
Let us fix a Hamiltonian structure of the form (2.6) and let $V$ be the Heisenberg vertex algebra of rank $N$ generated by states $b^{1},\dots,b^{N}$ satisfying
$\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N.$
The key observation is that we can identify the state space and the space of differential polynomials via the $\hbar$-linear morphism
@@
$\varphi$ is an isomorphism between vector spaces, and for any $f\in\mathcal{A}[\hbar]$, we have
@@
It follows from (2.8) that $\varphi$ induces an isomorphism between quotient spaces $\mathcal{F}[\hbar]$ and $V/TV$, and we will still denote this isomorphism by $\varphi$. For general vertex algebras, it is well-known that $V/TV$ admits a natural Lie algebra structure given by the 00-th product.
@@
(2.10) $[A,B]=\pi\left(\pi^{-1}(B)_{(0)}\pi^{-1}(A)\right),\quad A,B\in V/TV$
@@
Via $\varphi$, the above bracket on $V/TV$ can be identified to give a bracket on $\mathcal{F}[\hbar]$. Our first goal is to compute this bracket in terms of differential polynomials, and hence prove that this bracket gives a deformation quantization of the Hamiltonian structure (2.6). For this purpose, let us compute $\varphi(f)_{(n)}$ for $f\in\mathcal{A}$ and $n\geq 0$.
@@
where $V_{0}$ is the $\mathbb{C}$-vector space spanned by states of the form $b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$. Then it is obvious that for $a\in V_{0}$, we have $(\hbar^{m}a)_{(n)}=\hbar^{m}\cdot a_{(n)}$, and we can decompose
$a_{(n)}=\sum_{i\geq 0}a_{(n)}^{[i]},\quad a_{(n)}^{[i]}\colon V_{0}\to V_{i}.$
By using Corollary 2.2, if we denote
@@
then it follows from the commutation relation (2.5) that
$a_{(n)}^{[i]}=\sum_{(j_{1},\dots,j_{n})\in J^{i}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})},$
@@
Using Proposition 2.1, we have
@@
and by the above observation, we arrive at
@@
In order to obtain a non-vanishing summand, we must have $m_{i}\leq-1$, and therefore the only non-vanishing summand is given by the term with indices $m_{1}=\dots=m_{n}=-1$, namely,
@@
Remark 2.3.
This lemma also follows directly from the fact that $a_{(-1)}|0\rangle=a$ for any $a\in V$.
@@
By a similar argument as in the proof of Lemma 2.5, we see that $\varphi(f)_{(n)}^{[0]}=0$ for $n\geq 0$. Therefore, we only need to consider $\varphi(f)_{(n)}^{[m]}$ for $m\geq 1$. Without loss of generality, we may write
$\varphi(f)_{(n)}^{[m]}=\sum_{(j_{1},\dots,j_{\ell})\in J^{m}}C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{\ell}}_{(j_{\ell})},\quad C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}\in\mathbb{C}.$
It is more convenient to rewrite this operator into the form
@@
$b^{\beta_{1}}_{(-j_{1}-1)}\dots b^{\beta_{\ell-m}}_{(-j_{\ell-m}-1)},\quad j_{1},\dots,j_{\ell-m}\geq 0,$
and in particular, we may and do fix each operator $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ uniquely by requiring that it is symmetric with respect to indices $(\alpha_{1},k_{1}),\dots,(\alpha_{m},k_{m})$. Therefore, by using (2.9), we arrive at
@@
=\,= $\hbar^{m}m!\,s_{1}!\dots s_{m}!\,\eta^{\alpha_{1}\beta_{1}}\dots\eta^{\alpha_{m}\beta_{m}}A^{s_{1},\dots,s_{m}}_{\alpha_{1},\dots,\alpha_{m}}|0\rangle.$
Note that we have used the fact that $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ is symmetric with respect to indices. Hence, by using Lemma 2.5, we have
@@
hence we only need to compute
@@
By using the definition (2.7) of $\varphi$ and Proposition 2.1, we have
@@
Finally, we arrive at the desired formula by using the identity (2.8) and Corollary 2.6. The theorem is proved. ∎

### v3

@@ Acknowledgement
Part of the work was done during the JSPS International Research Fellowship of the author and was supported by JSPS KAKENHI Grant Number 23KF0114. He would like to thank Wenda Fang, Hiroshi Iritani, Si-Qi Liu, Youjin Zhang and Chunhui Zhou for very helpful discussions. He also thanks the anonymous referees for very helpful suggestions and comments to improve the presentation of the paper.
Data Availability
Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
@@ 2.1. Basics of vertex algebra
Let us review some basic facts about vertex algebras, and we refer to [9] for details. The paper [5] also serves as a reference for the topic.
A vertex algebra is given by the data $\left(V,|0\rangle,T,Y(-,z)\right)$, where $V$ is a vector space over $\mathbb{C}$ called the state space, $|0\rangle\in V$ is a special vector called the vacuum vector, $T\in\mathrm{End}(V)$ is a linear operator called the infinitesimal translation operator and $Y(-,z)\colon V\to\mathrm{End}(V)[[z^{\pm 1}]]$ is a linear morphism called the state-field correspondence. The state-field correspondence associates to each state $a\in V$ operators $a_{(n)}$ for all $n\in\mathbb{Z}$ given by
@@
We call $a_{(n)}$ the $n$-th mode of $a$. The data $\left(V,|0\rangle,T,Y(-,z)\right)$ of a vertex algebra are required to satisfy the following axioms:
@@
  • $T|0\rangle=0$ and for any $a\in V$, $\left[T,a_{(n)}\right]=-na_{(n-1)}$ and $Y(Ta,z)=\partial_{z}Y(a,z)$.
@@
$a_{\lambda}b]=\sum_{n\geq 0}\frac{\lambda^{n}}{n!}a_{(n)}b.$
It can be proved that the $\lambda$-brackets are polynomials in $\lambda$ with coefficients being states in $V$. They satisfy the following properties, where many useful identities for vertex algebras are encoded:
@@
    $Ta_{\lambda}b]=-\lambda[a_{\lambda}b]$, and $a_{\lambda}Tb]=(T+\lambda)[a_{\lambda}b]$.
@@
    $a_{\lambda}[b_{\mu}c]]=[b_{\mu}[a_{\lambda}c]]+[[a_{\lambda}b]_{\lambda+\mu}c]$.
For $a,b\in V$, we denote by :ab::ab: the normal order product, which is defined to be the state $a_{(-1)}b$. For general vertex algebras, the normal order product is neither commutative nor associative. It satisfies the following quasi-commutativity and quasi-associativity ([5]):
@@
$\int_{-T}^{0}d\lambda\,[a_{\lambda}b]=\sum_{n\geq 0}\left(\int_{-T}^{0}\frac{\lambda^{n}}{n!}d\lambda\right)\left(a_{(n)}b\right)=\sum_{n\geq 0}\frac{(-1)^{n}T^{n+1}}{(n+1)!}\left(a_{(n)}b\right).$
@@
(2.3) $a_{\lambda}:bc:]=:[a_{\lambda}b]c:+:b[a_{\lambda}c]:+\int_{0}^{\lambda}[[a_{\lambda}b]_{\mu}c]d\mu.$
Another standard fact about the normal order product is that
@@
In this paper, we will only consider the Heisenberg vertex algebra which we define now. Recall that a Heisenberg Lie algebra of rank $N$ is a $\mathbb{C}$-vector space spanned by $b^{\alpha}_{(n)}$ with $\alpha=1,\dots,N$ and $n\in\mathbb{Z}$, whose Lie bracket is specified by
(2.5) $\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N,\quad n,m\in\mathbb{Z},$
where $(\eta^{\alpha\beta})$ is a constant non-degenerate symmetric matrix and $\hbar$ is a formal parameter.
Remark 2.1.
Strictly speaking, the underlying space of the Heisenberg Lie algebra we define here is a free $\mathbb{C}[\hbar]$-module rather than a vector space. The only reason for the introduction of the parameter $\hbar$ is to make sense of classical limits by taking $\hbar\to 0$. In particular, all the definitions and axioms of vertex algebras make sense if the state space is a free $\mathbb{C}[\hbar]$-module and all endomorphisms are required to be $\hbar$-linear.
We proceed to define the Heisenberg vertex algebra. First define the state space $V$ to be the free $\mathbb{C}[\hbar]$-module spanned by states of the form
@@
In particular, we have the vacuum vector $|0\rangle\in V$, and we view $b^{\alpha}_{(n)}$ as endomorphisms on $V$ whose actions are determined by the commutation relation (2.5) and $b^{\alpha}_{(n)}|0\rangle=0$ for $n\geq 0$. As an example, we have
$b^{\alpha}_{(n)}\left(b^{\beta}_{(-k-1)}|0\rangle\right)=\begin{cases}b^{\alpha}_{(n)}b^{\beta}_{(-k-1)}|0\rangle,&n<0\\ \hbar\eta^{\alpha\beta}n\delta_{n-k-1,0}|0\rangle,&n\geq 0.\end{cases}$
Note that we have $b^{\alpha}_{(0)}=0$. We define the infinitesimal translation operator $T$ similarly by specifying its action, which is determined by $T|0\rangle=0$ and the commutation relation $\left[T,b^{\alpha}_{(n)}\right]=-nb^{\alpha}_{(n-1)}$. Finally, we define the state-field correspondence Y(,z)Y(-,z) by specifying the modes of each state. For a state $b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$, its modes are defined recursively by induction on $n$ [1]:
$|0\rangle_{(i)}$ $=\delta_{i,-1}\mathrm{Id}_{V},$
$(b^{\alpha}_{(i)}a)_{(j)}$ $=\sum_{\ell\geq 0}(-1)^{\ell}\binom{i}{\ell}\left(b^{\alpha}_{(i-\ell)}a_{(j+\ell)}-(-1)^{i}a_{(i+j-\ell)}b^{\alpha}_{(\ell)}\right),\quad i,j\in\mathbb{Z}.$
From now on, we will denote by $b^{\alpha}$ the state $b^{\alpha}_{(-1)}|0\rangle$, and it follows that
$Y(b^{\alpha},z)=\sum_{n\in\mathbb{Z}}b^{\alpha}_{(n)}z^{-n-1}.$
Lemma 2.1.
The date $(V,|0\rangle,T,Y(-,z))$ described above define a vertex algebra called the Heisenberg vertex algebra of rank $N$.
Proof.
We refer to Chapter 2 of [9] for a proof. ∎
For the Heisenberg vertex algebra, the modes of states admit particularly nice formulae in terms of the normal ordering which we shall define now. We call $b^{\alpha}_{(n)}$ an annihilator if $n\geq 0$ and a creator if $n<0$. Then, the normal ordering is defined for a sequence of modes of $b^{\alpha}$, which is denoted by $b^{\alpha_{1}}_{(k_{1})}\dots b^{\alpha_{n}}_{(k_{n})}:$, and its action is given by placing all annihilators on the right of all creators. It is important to note that the normal ordering is well-defined because creators commute with each other, so do annihilators. Now we proceed to give an explicit formula for the state-field correspondence of the Heisenberg vertex algebra.
@@
$Y\left(a,z\right)=(-1)^{\sum k_{i}}\sum_{m_{1},\dots,m_{n}\in\mathbb{Z}}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:z^{-n-\sum m_{i}}.$
@@
Let us prove by induction on $n$. The case $n=0$ trivially holds true. Assume that we have proved the proposition for some $n\geq 0$, and let us consider the state
$a=b^{\alpha}_{(-k-1)}b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V$
for some $k\geq 0$. First, it follows from $Y(T^{k}b^{\alpha},z)=\partial^{k}_{z}Y(b^{\alpha},z)$ that
@@
The field corresponding to $a$ is then computed by using the identity (2.4), and we arrive at the following expression for $Y(a,z)$:
$\begin{split}(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\in\mathbb{Z}\\ m<0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot b^{\alpha}_{(m-k)}\left(:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\right)z^{-n-\sum m_{i}-m-1}\end{split}$
$\begin{split}+(-1)^{k+\sum k_{i}}\sum_{\begin{subarray}{c}m_{1},\dots,m_{n}\in\mathbb{Z}\\ m\geq 0\end{subarray}}&\binom{m}{k}\binom{m_{1}}{k_{1}}\dots\binom{m_{n}}{k_{n}}\\ &\cdot\left(:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\right)b^{\alpha}_{(m-k)}z^{-n-\sum m_{i}-m-1}.\end{split}$
In the first summation, since we have mk<0m-k<0, it follows from the definition of the normal ordering that
$b^{\alpha}_{(m-k)}\left(:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\right)\,=\,:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:,\quad m<0.$
In the second summation, terms for $0\leq m\leq k-1$ vanish due to the fact that the binomial coefficient $\binom{m}{k}=0$, and for $m\geq k$ we have
$\left(:b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:\right)b^{\alpha}_{(m-k)}=:b^{\alpha}_{(m-k)}b^{\alpha_{1}}_{(m_{1}-k_{1})}\dots b^{\alpha_{n}}_{(m_{n}-k_{n})}:.$
@@
Using the above proposition as well as the definition of the normal ordering, we arrive at the following description of modes.
@@
Fix a state $a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$ for some $n\geq 0$ and $k_{i}\geq 0$, its modes are given by
$a_{(m)}=\sum_{\begin{subarray}{c}j_{1}\leq\dots\leq j_{n}\\ \beta_{1},\dots,\beta_{n}\end{subarray}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})}\in\mathrm{End}(V),\quad m\in\mathbb{Z},$
here $C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}$ are some integers depending on both the state $a$ and $m$, where the indices $\beta_{i}$ run over $1,\dots,N$ and $j_{i}$ run over $\mathbb{Z}$.
Finally, let us recall the conformal degree of the Heisenberg vertex algebra. For a state
$a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V,$
we call the integer
$\Delta_{a}=n+\sum_{i=1}^{n}k_{i}$
the conformal weight of $a$. Let us denote by $V^{\Delta}$ the subspace of $V$ consisting of homogeneous states of conformal weight $\Delta$, and we set $V^{\Delta}=0$ if $\Delta<0$. An endomorphism is called of conformal degree $d$ if it maps $V^{\Delta}$ to $V^{\Delta+d}$ for any $\Delta\geq 0$. For example, it is easy to see that the mode $b^{\alpha}_{(n)}$ is of conformal degree n-n for any $n\in\mathbb{Z}$.
Lemma 2.4.
If $a\in V^{\Delta}$, then $a_{(n)}$ is of conformal degree $\Delta-n-1$ for any $n\in\mathbb{Z}$. In particular, if $a\in V^{\Delta_{1}}$ and $b\in V^{\Delta_{2}}$, then $ab:\in V^{\Delta_{1}+\Delta_{2}}$.
Proof.
This is an immediate consequence of Proposition  2.2. ∎
@@ 2.2. Quantum Hamiltonian structure
In this section, we construct a deformation quantization of Hamiltonian structures of hydrodynamic types using the Heisenberg vertex algebra. To this end, we first recall the notion of a (classical) Hamiltonian structure, and we refer to [17] for an introduction to the general theory of Hamiltonian structures.
@@
$\mathcal{A}=\mathbb{C}\left[v^{\alpha,p}\colon\alpha=1,\dots,N;\,p\geq 0\right].$
From now on, we will fix such a local coordinate system. Define a derivation $\partial_{x}$ of $\mathcal{A}$ by
@@
Let us consider the following special Hamiltonian structure which is called the Hamiltonian structure of hydrodynamic type:
@@
where $(\eta^{\alpha\beta})$ is a non-degenerate symmetric constant matrix. It is proved [10] that the bracket (2.6) serves as a universal local model for Hamiltonian structures with hydrodynamic leading terms. Such Hamiltonian structures are ubiquitous in the theory of integrable hierarchies. For example, all Dubrovin-Zhang hierarchies [7] and all double ramification hierarchies [2] possess Hamiltonian structures of these kinds.
Let us fix a Hamiltonian structure of the form (2.6) and let $V$ be the state space of the Heisenberg vertex algebra of rank $N$ as described in Sect. 2.1. Recall that we have states $b^{1},\dots,b^{N}\in V$ whose modes satisfy the commutation relation
$\left[b^{\alpha}_{(n)},b^{\beta}_{(m)}\right]=\hbar\,\eta^{\alpha\beta}\,n\,\delta_{m+n,0},\quad\alpha,\beta=1,\dots,N,\quad n,m\in\mathbb{Z}.$
The key observation is that we can identify the state space $V$ and the space of differential polynomials $\mathcal{A}$ via the $\hbar$-linear map given by
@@
The map $\varphi$ defined in (2.7) is an isomorphism between vector spaces. For any $f\in\mathcal{A}[\hbar]$, we have
@@
It follows from (2.8) that $\varphi$ induces an isomorphism between quotient spaces $\mathcal{F}[\hbar]$ and $V/TV$, and we will still denote this isomorphism by $\varphi$. It is well-known that $V/TV$ admits a natural Lie algebra structure given by 0-th modes.
@@
(2.10) $A,B]=\pi\left(\pi^{-1}(B)_{(0)}\pi^{-1}(A)\right),\quad A,B\in V/TV$
@@
Via the isomorphism $\varphi$, the above bracket on $V/TV$ can be identified to give a Lie bracket on $\mathcal{F}[\hbar]$. Our first goal is to compute this bracket in terms of differential polynomials explicitly and hence to prove that this bracket gives a deformation quantization of the Hamiltonian structure (2.6). For this purpose, let us compute the modes $\varphi(f)_{(n)}$ for $f\in\mathcal{A}$ and $n\geq 0$.
@@
where $V_{0}$ is the $\mathbb{C}$-vector space spanned by states of the form $b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle$. Then it is clear that for $a=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{n}}_{(-k_{n}-1)}|0\rangle\in V_{0}$, we have $(\hbar^{m}a)_{(n)}=\hbar^{m}\cdot a_{(n)}$, and furthermore the modes of $a$ can be decomposed as
$a_{(n)}=\sum_{i\geq 0}a_{(n)}^{[i]},\quad a_{(n)}^{[i]}\colon V_{0}\to V_{i},\quad n\in\mathbb{Z}.$
It follows from Corollary  2.3 that these modes are of the forms
@@
here the coefficients $C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}$ are constant numbers independent of $\hbar$. Then the $\hbar$-dependence of modes is given by the commutation relation (2.5), hence we arrive at
(2.11) $a_{(n)}^{[i]}=\sum_{(j_{1},\dots,j_{n})\in J^{i}}C^{j_{1},\dots,j_{n}}_{\beta_{1},\dots,\beta_{n}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{n}}_{(j_{n})},$
@@
It follows from Proposition  2.2 that
@@
and by further using (2.11), we arrive at
@@
Since $b^{\alpha_{i}}_{(0)}=0$, it follows that the non-vanishing summand is with indices $m_{i}\leq-1$. Therefore, the only non-vanishing summand is given by the term with indices $m_{1}=\dots=m_{n}=-1$, namely,
@@
Proof.
This follows directly from the definition of $\varphi$. The corollary is proved. ∎
@@
By a similar argument as in the proof of Lemma  2.7, we see that $\varphi(f)_{(n)}^{[0]}=0$ for $n\geq 0$. Therefore, we only need to consider $\varphi(f)_{(n)}^{[m]}$ for $m\geq 1$. Without loss of generality, we simply assume
$\varphi(f)=b^{\alpha_{1}}_{(-k_{1}-1)}\dots b^{\alpha_{\ell}}_{(-k_{\ell}-1)}|0\rangle$
for some $\ell\geq 0$ and $k_{i}\geq 0$. Using (2.11), we may write
$\varphi(f)_{(n)}^{[m]}=\sum_{(j_{1},\dots,j_{\ell})\in J^{m}}C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}b^{\beta_{1}}_{(j_{1})}\dots b^{\beta_{\ell}}_{(j_{\ell})},\quad C^{j_{1},\dots,j_{\ell}}_{\beta_{1},\dots,\beta_{\ell}}\in\mathbb{Z}.$
It is more convenient to separate annihilators and creators in the above expression, and rewrite it into the form
@@
$b^{\beta_{1}}_{(-j_{1}-1)}\dots b^{\beta_{\ell-m}}_{(-j_{\ell-m}-1)},\quad j_{1},\dots,j_{\ell-m}\geq 0.$
Since annihilators commute with each other, we may and do fix each operator $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ uniquely by requiring that it is symmetric with respect to indices $(\alpha_{1},k_{1}),\dots,(\alpha_{m},k_{m})$. Therefore, by using (2.9), we arrive at
@@
$=\,$ $\hbar^{m}m!\,s_{1}!\dots s_{m}!\,\eta^{\alpha_{1}\beta_{1}}\dots\eta^{\alpha_{m}\beta_{m}}A^{s_{1},\dots,s_{m}}_{\alpha_{1},\dots,\alpha_{m}}|0\rangle.$
Note that we have used the fact that $A^{k_{1},\dots,k_{m}}_{\alpha_{1},\dots,\alpha_{m}}$ is symmetric with respect to indices. Hence, by using Lemma  2.7, we have
@@
hence to further simplify the above expression for $\varphi(f)_{(n)}^{[m]}\varphi(g)$ we only need to compute
@@
By using the definition (2.7) of $\varphi$ and Proposition  2.2, we have
@@
Finally, we arrive at the desired formula by using the identity (2.8) and Corollary  2.8. The theorem is proved. ∎